Let A =

(8 -1 2

2 0 -5)


B =

(-1 7

3 -2

1 5)


and C =

(2 1

3 5)


(a) Calculate AB and A + B if they exist.

(b) Verify that (AB)C = A(BC).

(c) Calculate C-1 A.


1
Expert's answer
2021-11-09T16:02:03-0500

a)

(A+B) does not exist, since matrices are of different sizes (2x3 and 3x2)

"AB=\begin{pmatrix}
 8 & -1 &2\\
 2 & 0&-5
\end{pmatrix}\begin{pmatrix}
 -1 & 7 \\
 3 & -2\\
1&5
\end{pmatrix}=\begin{pmatrix}
 -8-3+2 & 56+2+10 \\
 -2-5 & 14-25
\end{pmatrix}=\begin{pmatrix}
 -9 & 68 \\
 -7 & -11
\end{pmatrix}"


b)

"(AB)C=\begin{pmatrix}
 -9 & 68 \\
 -7 & -11
\end{pmatrix}\begin{pmatrix}
 2 & 1 \\
 3 & 5
\end{pmatrix}=\begin{pmatrix}
 -18+204 & -9+340 \\
 -14-33 & -7-55
\end{pmatrix}=\begin{pmatrix}
 186 & 331 \\
 -47 & -62
\end{pmatrix}"


"BC=\begin{pmatrix}
 -1 & 7 \\
 3 & -2\\
1&5
\end{pmatrix}\begin{pmatrix}
 2 & 1 \\
 3 & 5
\end{pmatrix}=\begin{pmatrix}
 -2+21 & -1+35 \\
 6-6 & 3-10\\
2+15&1+25
\end{pmatrix}=\begin{pmatrix}
 19 & 34 \\
 0 & -7\\
17&26
\end{pmatrix}"


"A(BC)=\begin{pmatrix}
 8 & -1 &2\\
 2 & 0&-5
\end{pmatrix}\begin{pmatrix}
 19 & 34 \\
 0 & -7\\
17&26
\end{pmatrix}=\begin{pmatrix}
 152+34 & 272+7+52 \\
 38-85 & 68-130
\end{pmatrix}="


"=\begin{pmatrix}
 186 & 331 \\
 -47 & -62
\end{pmatrix}"


c)

for C-1:

"C_*=\begin{pmatrix}
 5 & -3 \\
 -1 & 2
\end{pmatrix}" , "C_*^T=\begin{pmatrix}
 5 & -1 \\
 -3 & 2
\end{pmatrix}"


"|C|=10-3=7"


"C^{-1}=\frac{C_*^T}{|C|}=\frac{}{}=\frac{1}{7}\begin{pmatrix}
 5 & -1 \\
 -3 & 2
\end{pmatrix}"


"C^{-1}A=\frac{1}{7}\begin{pmatrix}
 5 & -1 \\
 -3 & 2
\end{pmatrix}\begin{pmatrix}
 8 & -1 &2\\
 2 & 0&-5
\end{pmatrix}=\frac{1}{7}\begin{pmatrix}
 40-2 & -5 &10+5\\
 -24+4 & 3&-6-10
\end{pmatrix}="


"=\frac{1}{7}\begin{pmatrix}
 38 & -5 &15\\
 -20 & 3&-16
\end{pmatrix}"